# Multi-agent cooperation through learning-aware policy gradients

### v1

@@ 7 Conclusion
We have shown that learning awareness allows reaching high returns in challenging social dilemmas, designed to make independent learning difficult. We identified two key conditions for this to occur. First, we found it necessary to take into account the stochastic minibatched nature of the updates used by other agents. This is one distinguishing aspect of the COALA-PG learning rule proposed here, which translates into a significant performance advantage over prior methods. Second, learning-aware agents had to be embedded in a heterogeneous group containing non-learning-aware agents. An interesting question for future work is whether this condition can be generalized to mixtures of agents that learn at multiple time scales, beyond the all-or-none case considered here.
@@
Acknowledgements.
We are grateful to Joel Z. Leibo and Sasha Vezhnevets for an insightful discussion, and to Rif A. Saurous for providing detailed comments on a version of this manuscript.
Appendix A Experimental details
A.1 Environments
A.1.1 Iterated prisoner’s dilemma (IPD)
@@ A.1.2 CleanUp-lite
  • State: The world is a grid of size 3, arranged horizontally. The right extremity is the river, and the left one the orchard. The world state also contains the pollution level of the river $P$, number of apples currently available in the orchard $A$, the position of each agent, and their respective zapped state
@@
  • Dynamic: the environment evolves at every timestep in the following order:
@@
      The pollution level increases by 1 with a fixed probability $p_{\text{pollution}}=0.5$, and capped at $P_{\text{max}}=5$.
@@
      The apple grows by 1 with probability $1-\min(1,P/P_{\text{threshold}})$, where $P_{\text{threshold}}=3$, and capped at $A_{\text{max}}=4$.
@@
      Then, agent in the "orchard" cell who is not zapped, has a probability $A/A_{\text{max}}$ of picking up an apple, independently from one another, in which case the apple level decreases by 1 (capped at 0).
@@
      An agent in the "river" cell who is not zapped has a probability $p_{\text{clean}}=0.7$ of decreasing the pollution level by 1, independently of other agents (capped at 0).
@@
      Finally, an agent zapping has a $p_{\text{zap}}=0.8$ probability of successfully zapping the opponent who is in the same cell. If the zapping is successful, the opponent is frozen for $t_{\text{zap}}=3$ timesteps, during which it is frozen and cannot be further zapped.
  • Initial state: Agents are randomly placed on the grid, unzapped, $A$ is set to 00, $P$ to $P_{\text{threshold}}$.
  • Observation: the observation is fully observable. Each agent sees the position of each agent encoded as concatenated one-hot vectors indicating the position in the grid, the normalized number of apples $A/A_{\text{max}}$, normalized pollution level $P/P_{\text{max}}$, the normalized remaining timeout status (zapped or not), and if so, a countdown indicating how many timestep is of each agent ($t_{\text{agent}}/t_{\text{zap}}$ where $t_{\text{agent}}$ is the time left before the agent unfreezes). The observation is symmetric.
@@ A.2.1 Meta agents
    First, a meta batch of opponents is sampled. Each opponent is hierarchically sampled by first determining whether it is a naive opponent (with probability $p_{\text{naive}}$), and then sampling uniformly from $\{\phi_{i}\}$ or $\{\psi_{i}\}$ accordingly. The sampling is done with replacement, and disallow sampling of oneself.
@@
    For each opponent, generate a batch of $B$ trajectories of length $T*M$, where $M$ is the number of inner episode, and $T$ the episode length of the environment. Crucially, after every $T$ steps, the environment terminates and is reset, and, if the opponent is naive, the previous batch of length $T$ trajectory is used to update its parameter following an update rule RL of choice.
@@
    For each collected batched trajectories, the policy gradient of the meta agent parameter is computed following the COALA-PG update rule (or other baselines, c.f. E) if the opponent is naive, and the standard policy gradient otherwise otherwise (i.e. the batch and meta batch dimensions are flattened). Crucially, the done signal from the inner episodes are ignored. The gradient is then averaged, and the parameter updated.
A.2.2 Naive agent
@@
The naive agents parameters are used to initialize the naive opponents when training the meta agents, but the resulting "trained" parameters is discarded. The initialization may or may not be nonetheless updated during training. In more challenging environment however, training from scratch until good performance is achieved in a single meta trajectory may require prohibitely many inner episodes. To avoid this, in some of our experiments, at each training iteration, we set each $\{\psi_{i}\}$ to be equal to one of the $\{\phi_{i}\}$. This ensures that naive agents are initialized as an already capable agent, and is possible due to our choice of common architecture between naive and meta agents (c.f. below). In that case, we say that the naive agents are dynamic. Otherwise, naive agent is always initailized at one of a predefined static set of parameters.
A.2.3 Reinforcement learning
We experiment with both Advantage Actor-Critic (A2C) (Mnih et al., 2016) and Proximal Policy Optimization (PPO) Schulman et al. (2017) on our COALA policy gradient estimate. We details here how to adapt our policy gradient to these methods, together with additional tricks.
We remind for reference, that by denoting by $m_{l}$ the inner episode index corresponding to the meta episode time step $l$, the COALA Policy gradient can be rewritten as
We remind that the standard policy gradient, which would be used to train against e.g. another meta agent, is
Value function estimation.
Value functions are used for reducing the variance in the policy gradient estimator, by typically subtracting a baseline from the return estimator. In the above equation 7, the straightforward value function function to learn is
On expectation, the second term equals can be simplified, and the expression then becomes
Which has an additional $\frac{1}{B}$ factor on the left term compared to a value function that would need to be learnt when playing e.g. against a meta agent. This is undesirable for several reasons, one of which being that as $B$ increases, the value target becomes increasingly insensitive to the inner episode return, which makes learning difficult. Another reason is that the target magnitude between when playing against a naive agent or a meta agent can significantly differ.
We circumvent this by instead always learning the batch-unaware value function, i.e.
As such, the same target is used for both when playing against a naive or a meta agent.
In practice, one can use the monte carlo estimates to learn it (or with variance reduction techniques such as td-lambda). Optionally, the second term can be replaced by the return over future inner episodes averaged over inner batch, for meta-episodes against naive learners.
We now see how the above value estimation can be used to update the policy following COALA-PG.
Policy update.
Ultimately we want an unbiased estimate of the advantage function, since this would allow the usage of algorithms like PPO or A2C.
We can reformulate the expression using $\hat{V}$ as follows:
A candidate is thus to simply take the monte carlo estimate of the above.
Additional tricks.
Here, we list a few tweaks of classical reinforcement learning trick for our application.
  • Advantage normalization: as is common with PPO implementation, we investigate the use of advantage normalization. Given a batched trajectory of advantage estimation over which the policy should be updated, the trick consists in centering the advantage estimates over the batched trajectory. Empirically, we found out that when playing against a mixture of naive and meta learners, it was beneficial to apply the centering separately for the 2 types of meta-trajectories.
  • Value normalization: Given that meta trajectories can be extremely long, and thus the return of very high variance, we investigate the use of value normalization, a crude approximation of the PopArt layer (Hessel et al., 2019). Effectively, for each batched trajectory on which the value loss is to be computed, we first normalize the empirical return by centering it to unit variance and zero mean, and regress the value estimate against it. Then, for the value prediction used for the policy update, the prediction is de-normalized (using the statistics from the batch). This allows for improved learning of value estimate with deep learning, at the cost of introducing non stationarity.
A.3 Architecture
@@
A.4 Hyperparameters for each experiment
@@
population_size (meta) $1$ $4$
population_size (naive) $10$ $4$
@@
value_normalization $\{False,True\}$$\{False,True\}$
[…]
value_discount ($\gamma$) $\{0.999,1.0\}$$\{0.99,1.0\}$
gae_lambda ($\lambda_{\text{gae}}$) $\{0.98,1.0\}$$\{1.0\}$
@@
Coala Batch Unaware M-FOS Coala Batch Unaware M-FOS
algorithm PPO PPO PPO A2C A2C A2C
@@
clip_value True True True True True True
entropy_regularization 00 00 00 00 00 00
advantage_normalization False True True True True True
value_normalization True True True False False False
reward_rescaling $1$ $1$ $1$ $0.1$ $0.1$ $0.1$
$\gamma$ $1$ $0.98$ $0.999$ $1$ $1$ $1$
$\lambda_{\text{td}}$ $1$ $1$ $1$ $1$ $1$ $1$
$\lambda_{\text{gae}}$ $1$ $1$ $1$ $1$ $1$ $1$
optimizer ADAM ADAM ADAM SGD SGD SGD
@@
learning_rate $0.003$ $0.001$ $0.001$ $0.03$ $0.03$ $0.1$
max_grad_norm $1$ $1$ $1$ $1$ $1$ $1$
@@
value_normalization False False False False
reward_rescaling $0.05$ $0.05$ $1$ $0.1$
value_discount ($\gamma$) $0.99$ $0.99$ $0.99$ $0.99$
@@
Appendix B The analytical iterated prisoner’s dilemma
For the experiments in Section 4 and 5, we analytically compute the discounted expected return of an infinitely iterated prisoner’s dilemma, and its parameter gradients. Automatic differentiation allows us then to explicitly backpropagate through the learning trajectory of naive learners, to compute the ground-truth meta update. In the following, we provide details on this approach.
@@ Appendix B The analytical iterated prisoner’s dilemma
$M=\begin{bmatrix}\sigma(\phi^{1}_{1:4})\odot\sigma(\phi^{2}_{1:4}),\sigma(\phi^{1}_{1:4})\odot(1-\sigma(\phi^{2}_{1:4})),(1-\sigma(\phi^{1}_{1:4}))\odot\sigma(\phi^{2}_{1:4}),(1-\sigma(\phi^{1}_{1:4}))\odot(1-\sigma(\phi^{2}_{1:4}))\end{bmatrix}^{\top}$
@@
For the LOLA experiments of Section 5, we compute the ground-truth LOLA-DICE updates equation 5 by initializing a naive learner with the opponent’s parameters, simulate $M$ naive updates (look-aheads) following partial derivatives of $J^{i}$, and backpropagating through the final return $J^{i}\left(\phi^{i},\phi^{-i}+\sum_{q=1}^{M}\Delta_{q}\phi^{-i}\right)$, including backpropagating through the learning trajectory. For Figure 4, we train two separate LOLA agents against each other and report the training curves of the first agent (the training curves of the second agent are similar, data not shown). Using self-play instead of other-play resulted in similar results with the same main conclusions (data not shown). For all experiments, we used the following hyperparemeters: $\gamma=0.95$, $\eta_{\text{meta}}=0.005$, $\eta_{\text{naive}}=5$ (except for 1-look-ahead, where we used $\eta_{\text{naive}}=10$). For Figure 4C, we used a convex mixture of the LOLA-DICE gradient and the partial gradient on $J^{i}(\phi^{i},\phi^{-i})$ with mixing factor $p_{\text{naive}}$. We used $p_{\text{naive}}=1,1,0.75,0.6,0.4$ for look-aheads 1, 2, 3, 10 and 20 respectively.
Appendix C Proofs
Theorem C.1.

### v2

@@ 6 Conclusion
We have shown that learning awareness allows reaching high returns in challenging social dilemmas, designed to make independent learning difficult. We identified two key conditions for this to occur. First, we found it necessary to take into account the stochastic minibatched nature of the updates used by other agents. This is one distinguishing aspect of the COALA-PG learning rule proposed here, which translates into a significant performance advantage over prior methods. Second, learning-aware agents had to be embedded in a heterogeneous group containing non-learning-aware agents.
@@
Acknowledgements
We would like to thank Maximilian Schlegel, Yanick Schimpf, Rif A. Saurous, Joel Leibo, Alexander Sasha Vezhnevets, Aaron Courville, Juan Duque, Milad Aghajohari, Razvan Ciuca, Gauthier Gidel, James Evans and the Google Paradigms of Intelligence team for feedback and enlightening discussions. GL and BR acknowledge support from the CIFAR chair program. EE acknowledges support from a Vanier scholarship from the government of Canada.
Appendix A A2C and PPO implementations of COALA-PG
We use both Advantage Actor-Critic (A2C) (Mnih et al., 2016) and Proximal Policy Optimization (PPO) Schulman et al. (2017) for our COALA policy gradient estimate. We detail here how to merge these methods with our COALA-PG method.
A.1 REINFORCE estimator
For the reader’s convenience, we display the COALA policy gradient below. We remind for reference, that $m_{l}$ the inner episode index corresponding to the meta episode time step $l$.
The batch-unaware COALA policy gradient, which we use as a baseline method for shaping naive learners, is given by
Note that when we play against other meta agents instead of naive learners, all parallel POMDP trajectories in the batch are independent, and hence we can correctly use the batch-unaware COALA policy gradient for this setting.
Finally, the M-FOS policy gradient (c.f. Appendix G) is given by
$\nabla_{\phi^{i}}\bar{J}(\phi^{i})$ $=\mathbb{E}_{\bar{P}^{\phi^{i}}}\left[\sum_{b=1}^{B}\sum_{l=1}^{MT}\nabla_{\phi^{i}}\log\pi^{i}(a^{i,b}_{l}\mid h^{i,b}_{l})\left(\sum_{k=l}^{Tm_{l}}R_{k}^{i,b}+\frac{1}{B}\sum_{b^{\prime}=1}^{B}\sum_{k=Tm_{l}+1}^{MT}R_{k}^{i,b^{\prime}}\right)\right].$ (7)
The difference between M-FOS and COALA-PG is the $\frac{1}{B}$ scaling factor for the current inner episode return. This scaling factor is crucial for a correct balance between gradient contributions arising from the current inner episode, and future inner episodes. Without this scaling factor, the contributions from future inner episodes required for learning to shape the co-players vanish for large inner batch sizes.
We can construct REINFORCE estimators by sampling directly from the above expectations. However, this leads to policy gradients with prohibitively large variance. Hence, in the following sections we will derive improved advantage estimators to reduce the variance of the policy gradient estimates.
A.2 Value function estimation
One of the easiest ways to use value functions for reducing the variance in the policy gradient estimator, is to subtract a baseline from the return estimator. In the COALA-PG equation 5, the straightforward value function to learn is
As the environment is reset after each inner episode, the second term can be simplified by merging expectations over the different parallel trajectories:
Which has an additional $\frac{1}{B}$ factor on the left term compared to a conventional value function that would need to be learnt when playing e.g. against another meta agent equation 6. This is undesirable for several reasons, one of which being that as $B$ increases, the value target becomes increasingly insensitive to the inner episode return, which makes learning difficult. Another reason is that the target magnitude between when playing against a naive agent or a meta agent can significantly differ. Finally, for simplicity reasons, we want a value function that we can use both when playing against naive learners, as well as other meta agents.
We can solve these issues by instead learning the value function for the batch-unaware returns, and introducing some specialized reweighing when playing against naive learners, which we will see later.
As such, the same value function can be used for both when playing against a naive or a meta agent. In practice, we trade off variance and bias for learning the value function by using TD($\lambda$) targets. Algorithm 1 shows how to compute such targets with a general algorithm, which we later can also repurpose for Generalized Advantage Estimation (Schulman et al., 2016) and M-FOS value functions. For computing the TD($\lambda$) targets for learning our value functions, we use normalize_current_episode=False and average_future_episodes=False when the given trajectory batch originates from playing against another meta agent.
Input: $r_{t}$, $discount$, $v_{t}$, $\lambda$, $average\_future\_episodes$, $normalize\_current\_episode$, $inner\_episode\_length$
Output: returns
$seq\_len\leftarrow r_{t}.shape[1]$
$batch\_size\leftarrow r_{t}.shape[0]$
if normalize_current_episode then
      $normalization\leftarrow batch\_size$
else
      $normalization\leftarrow 1$
$episode\_end\leftarrow(\mathbf{range}(seq\_len)\mod inner\_episode\_length)==(inner\_episode\_length-1)$
$acc\leftarrow v_{t}[:,-1]$
$global\_acc\leftarrow mean(v_{t}[:,-1])$
for $t=seq\_len-1$ to 00 do
       if $average\_future\_episodes$ and $episode\_end[t]$ then
             $acc\leftarrow global\_acc$
      $acc\leftarrow r_{t}[:,t]/normalization+discount\times((1-\lambda)\times v_{t}[:,t]+\lambda\times acc)$
       $global\_acc\leftarrow mean(r_{t}[:,t]+discount\times((1-\lambda)\times v_{t}[:,t]+\lambda\times global\_acc))$
       $returns[:,t]\leftarrow acc$
return returns
Algorithm 1 Batch Lambda Returns
A.3 Generalized Advantage Estimation
We now see how the above value estimation can be used to update the policy following COALA-PG. Ultimately we want an unbiased estimate of the advantage function, as this allows the usage of algorithms like PPO or A2C.
We can reformulate the expression using $\hat{V}$ as follows:
A simple advantage estimator would be the Monte-Carlo estimate of the above. However, we can trade-off variance with bias by using the Generalized Advantage Estimator (Schulman et al., 2016). Using similar logic as for the equation above, we can compute the COALA version of the GAE by reusing the batched_lambda_returns algorithm (c.f. Algorithm 1 as follows:
  • Instead of the rewards of the trajectory, we provide the TD errors $\delta_{t}=r_{t}+\gamma\hat{V}_{t+1}-\hat{V}_{t}$ as input for r_t.
  • We provide $\gamma\lambda$ as input for discount
  • We provide $1.0$ as input for $\lambda$
  • We put average_future_episodes and normalize_current_episode both on True.
For computing the GAE for the batch-unaware COALA-PG baseline, we follow the same approach except putting average_future_episodes and normalize_current_episode both on False. For computing the GAE for the M-FOS baseline (c.f. G), we follow the same approach except putting average_future_episodes on True and normalize_current_episode on False.
A.4 A2C and PPO implementations
We can now use the above advantage estimates directly into A2C and PPO implementations. Below, we list a few tweaks of classical reinforcement learning tricks that we used in our implementation.
  • Advantage normalization: as is common with PPO implementations, we investigate the use of advantage normalization. Given a batched trajectory of advantage estimation over which the policy should be updated, the trick consists in centering the advantage estimates over the batched trajectory. Empirically, we found out that when playing against a mixture of naive and meta learners, it was beneficial to apply the centering separately for the 2 types of meta-trajectories (playing against naive learners or playing against other meta agents).
Appendix B Experimental details
B.1 Environments
B.1.1 Iterated prisoner’s dilemma (IPD)
@@ B.1.2 CleanUp-lite
  • State: The world is a 2D grid of size $5\times 4$. The right column is the river, and the left one the orchard. Cells in the river column can be occupied by dirt. Cells in the orchard column can be occupied by an apple. The world state also contains the position of each agent, and their respective zapped state
@@
  • Dynamics: the environment evolves at every timestep in the following order:
@@
      When there is at least one cell in the river column that is not occupied by dirt, a new patch of dirt is spawned with probability $p_{\text{pollution}}=0.35$, and placed randomly on one of the free cells in the river column.
@@
      When there is at least one cell in the orchard column that is not occupied by dirt, a new apple is spawned with probability $p_{\text{apple}}=1-\min(1,P/P_{\text{threshold}})$, where $P_{\text{threshold}}=3$ and $P$ the total number of dirt cells in the environment. The spawned apple is placed randomly on one of the free cells in the orchard column.
@@
      When an agent that is not zapped visits a cell with an apple, it harvests the apple and gets a reward of 1. The apple is replaced by an empty cell
@@
      When an agent that is not zapped visits a cell with a dirt patch, it cleans the dirt patch and replaces it by an empty cell.
@@
      Finally, an agent zapping has a $p_{\text{zap}}=0.9$ probability of successfully zapping the co-player, if the co-player is maximally 2 cells away from the agent. If the zapping is successful, the opponent is frozen for $t_{\text{zap}}=5$ timesteps, during which it is frozen and cannot be further zapped.
    6. 6.
[…]
  • Initial state: Agents are randomly placed on the grid, unzapped, there are no apples at initialization and 3 dirt patches randomly placed in the river column.
  • Observation: the observation contains full information about the environment. Each agent sees the position of each agent encoded as flattened one-hot grid indicating the position in the grid, the full grid as a flattened grid with one-hot objects (apple, dirt, empty), and the state of all agents (zapped or non-zapped). The observation is symmetric.
@@ B.2.1 Meta agents
    First, a meta batch of opponents is sampled. Each opponent is hierarchically sampled by first determining whether it is a naive opponent (with probability $p_{\text{naive}}$), and then sampling uniformly from $\{\phi_{i}\}$ or $\{\psi_{i}\}$ accordingly. The sampling is done with replacement, and disallowing sampling of oneself.
@@
    For each opponent, generate a batch of $B$ trajectories of length $TM$, where $M$ is the number of inner episode, and $T$ the episode length of the environment. Crucially, after every $T$ steps, the environment terminates and is reset, and, if the opponent is naive, the previous batch of length $T$ trajectories is used to update its parameter following a RL update rule of choice.
@@
    For each collected batched trajectories, the policy gradient of the meta agent parameter is computed following the COALA-PG update rule (or other baselines, c.f. G) if the opponent is naive, and the standard policy gradient otherwise otherwise (i.e. the batch and meta batch dimensions are flattened). Crucially, the done signals from the inner episodes are ignored. The gradient is then averaged, and the parameter updated.
B.2.2 Naive agent
@@
The naive agent parameters are used to initialize the naive opponents when training the meta agents, but the resulting trained parameters are discarded. The initialization may or may not be nonetheless updated during training. In a more challenging environment however, training from scratch until good performance is achieved in a single meta trajectory may require prohibitively many inner episodes. To avoid this, in some of our experiments, at each training iteration, we set each $\{\psi_{i}\}$ to be equal to one of the $\{\phi_{i}\}$. This ensures that naive agents are initialized as an already capable agent, and is possible due to our choice of common architecture between naive and meta agents (c.f. below). In that case, we say that the naive agents are dynamic. Otherwise, naive agent is always initailized at one of a predefined static set of parameters.
B.3 Architecture
@@
B.4 Hyperparameters for each experiment
@@
population_size (meta) $1$ $3$
population_size (naive) $10$ $3$
@@
value_discount ($\gamma$) $\{0.999,1.0\}$$\{1.0\}$
@@
Coala Batch Unaware M-FOS LOLA Coala Batch Unaware M-FOS LOLA
algorithm PPO PPO PPO - A2C A2C A2C -
@@
clip_value True True True - True True True True
entropy_regularization 00 00 00 00 00 00 00 00
advantage_normalization True True True True True True True True
$\gamma$ $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$
$\lambda_{\text{td}}$ $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$
reward_rescaling $0.1$ $1$ $1$ $1$ $0.1$ $0.1$ $0.1$ $0.1$
$\lambda_{\text{gae}}$ $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$
optimizer ADAM ADAM ADAM SGD SGD SGD SGD SGD
@@
learning_rate $0.001$ $0.001$ $0.003$ $0.1$ $0.1$ $0.03$ $0.03$ $0.03$
max_grad_norm $1$ $1$ $1$ - $1$ $1$ $1$ $1$
@@
reward_rescaling $0.05$ $0.05$ $0.1$ $0.1$
value_discount ($\gamma$) $0.99$ $0.99$ $0.99$ $1$
@@
Appendix C The analytical iterated prisoner’s dilemma
For the experiments in Section 4 and 4.3, we analytically compute the discounted expected return of an infinitely iterated prisoner’s dilemma, and its parameter gradients. Automatic differentiation allows us then to explicitly backpropagate through the learning trajectory of naive learners, to compute the ground-truth meta update. In the following, we provide details on this approach.
@@ Appendix C The analytical iterated prisoner’s dilemma
$M$ $=$
$\begin{bmatrix}\sigma(\phi^{1}_{1:4})\odot\sigma(\phi^{2}_{1:4}),\sigma(\phi^{1}_{1:4})\odot(1-\sigma(\phi^{2}_{1:4})),(1-\sigma(\phi^{1}_{1:4}))\odot\sigma(\phi^{2}_{1:4}),(1-\sigma(\phi^{1}_{1:4}))\odot(1-\sigma(\phi^{2}_{1:4}))\end{bmatrix}^{\top}$
@@
For the LOLA experiments of Section 4.3, we compute the ground-truth LOLA-DICE updates equation 4 by initializing a naive learner with the opponent’s parameters, simulate $M$ naive updates (look-aheads) following partial derivatives of $J^{i}$, and backpropagating through the final return $J^{i}\left(\phi^{i},\phi^{-i}+\sum_{q=1}^{M}\Delta_{q}\phi^{-i}\right)$, including backpropagating through the learning trajectory. For Fig. 4, we train two separate LOLA agents against each other and report the training curves of the first agent (the training curves of the second agent are similar, data not shown). Using self-play instead of other-play resulted in similar results with the same main conclusions (data not shown). For all experiments, we used the following hyperparemeters: $\gamma=0.95$, $\eta_{\text{meta}}=0.005$, $\eta_{\text{naive}}=5$ (except for 1-look-ahead, where we used $\eta_{\text{naive}}=10$). For Figure 4C, we used a convex mixture of the LOLA-DICE gradient and the partial gradient on $J^{i}(\phi^{i},\phi^{-i})$ with mixing factor $p_{\text{naive}}$. We used $p_{\text{naive}}=1,1,0.75,0.6,0.4$ for look-aheads 1, 2, 3, 10 and 20 respectively.
C.1 Additional results on the analytical IPD
Learning-aware agents extort naive learners following Zero-Determinant-like extortion strategies.
Figure 3A shows that learning-aware agents trained against naive learners find a policy that extorts the naive learners into unfair cooperation. Here, we investigate the resulting extortion policies in more detail, and show that they are similar to the Zero-Determinant extortion strategies discovered by Press and Dyson (2012). Zero-determinant extortion strategies are parameterized by $\chi$ and $\phi$ as follows (with $(T,R,P,S)=(2,1,0,-1)$ the rewards of the prisoner’s dilemma):
$\begin{split}p_{1}&=1-\phi(\chi-1)\frac{R-P}{P-S}\\ p_{2}&=1-\phi\left(1+\chi\frac{T-P}{P-S}\right)\\ p_{3}&=\phi\left(\chi+\frac{T-P}{P-S}\right)\\ p_{4}&=0\end{split}$ (18)
with $\chi\geq 1$ and $0<\phi\leq\frac{P-S}{(P-S)+\chi(T-P)}$. For $\chi=1$ and $\phi=\frac{P-S}{(P-S)+(T-P)}$ we recover the tit-for-tat strategy, representing the fair shaping strategy, whereas for higher values of $xi$, the resulting policies extort the naive learner into unfair cooperation. Note that Press and Dyson (2012) did not consider a $p_{0}$ parameter, as there theory is independent of the choice for $p_{0}$.
To investigate whether our learned co-player shaping policies are related to ZD extortion strategies, we take the converged policy $\sigma(\phi^{i})$ after training with the pure shaping objective (c.f. Figure 3A), and fit the parameters $(\chi,\phi)$ to the regression loss $\|\sigma(\phi^{i})[1:5]-p^{ZD}(\chi,\phi)\|^{2}$, with $p^{ZD}(\chi,\phi)$ the ZD extortion policy of Eq. 18. Figure 8 shows that policies learned with the co-player shaping objective can be well aproximated by ZD extortion policies, whereas random policies cannot. The ZD extortion policies of Eq. 18 consider undiscounted infinitely repeated matrix games, whereas we consider discounted infinitely repeated prisoner’s dilemma with discount $\gamma=0.999$. Furthermore, our shaping objective considers the cumuluted returns over the whole learning trajectory of the naive learner, in contrast to ZD extortion strategies that are optimized for the maximizing the return of the last inner episode. Hence, we should not expect an exact match between the learned policies $\sigma(\phi^{i})$ and the ZD extortion strategies.
Figure 8: Histogram of the regression losses after fitting the $(\chi,\phi)$ parameters to the learned co-player shaping policies from Figure 3A for 64 random seeds, versus fitting the $(\chi,\phi)$ parameters to 64 uniform random policies.
Mutual unconditional defection is not a Nash equilibrium in the mixed group setting.
First, we check numerically whether mutual unconditional defection results in a zero gradient, a necessary condition for being a Nash equilibrium. As a zero probability corresponds to infinite logits, we parameterize our policy now directly in the probability space instead of logit space, and consider projected gradient ascent to the probability simplex, i.e. clipping the updated parameters between 0 and 1. For non-zero mixing factors $p_{\text{naive}}$, this results in a projected gradient that is 0 everywhere, except for the parameter corresponding to the DC state. For shaping naive learners, it is beneficial to reward co-players that cooperate by also cooperating with non-zero probability afterwards. Hence, when an agent with a pure defection policy plays against naive learners, the resulting gradient will push it out of the pure defection policy.
Figure 9A shows that when we train unconditional defection policies in the mixed group setting with the same hyperparameters as for Figure 3C, the agents escape mutual defection and learn to cooperate. Note that the agents quickly learn how to shape naive learners, and that it takes a bit longer to learn full cooperation, as the shaping objective does not provide pressure to increase the cooperation probability in the starting state. However, as the shaping policies of the meta agents are not any longer unconditional defection, playing against other meta agents provides a pressure to increase the cooperation probability, eventually leading to a phase transition towards cooperation. Figure 9B shows the parameter trajectory in logit space over training, showing that indeed the agents adjust quickly their parameters for shaping, and eventually also the initial state cooperation probability, leading to cooperation against other meta agents. As our policies are parameterized in logit space, we initialize them to $\log 0.01$ instead of exactly to zero cooperation probability to avoid infinities.
Figure 9: (A) Average reward during training in a mixed group setting with both agents starting from an unconditional defection policy, when evaluating the learned policy versus naive agents (shaping reward) and versus the other learned policy (other-play reward). (B) Parameter trajectory in logit space of the first agent (the second agent has similar learning trajectories, data not reported). Shaded regions indicate 0.25 and 0.75 quantiles, and solid lines the median over 8 random seeds.
Tit for Tat is not a Nash equilibrium in the mixed group setting.
Figure 10 repeats the same analysis but now starting from Tit for Tat policies, showing that mutual Tit for Tat is not a Nash equilibrium in our mixed pool setting. As we show in Figure 10C, this is caused by the possibility to shape naive learners faster by deviating from a strict tit for tat policy. Note that even though the resulting policies are not perfect tit for tat, they still fully cooperate when played against each other.
Figure 10: (A) Average reward during training in a mixed group setting with both agents starting from a tit for tat policy, when evaluating the learned policy versus naive agents (shaping reward) and versus the other learned policy (other-play reward). (B) Parameter trajectory in logit space of the first agent (the second agent has similar learning trajectories, data not reported). (C) The reward of a main agent when playing against a naive learner over a trajectory of 20 naive learning steps, showing that the policy learned after convergence in the mixed group setting shapes a naive learner faster compared to a tit for tat policy. Shaded regions indicate 0.25 and 0.75 quantiles, and solid lines the median over 8 random seeds.
Appendix D Proofs
Theorem D.1.